find the LCM of 8,12,and72

Answer by stanbon(48510)

find the LCM of 8,12,and72
---
8 = 2^3
12 = 2^2*3
72 = 2^3*3^2
------------------
The lcm must have each of the prime factors at their highest power:
lcm = 2^3*3^2 = 72
--------------------
Cheers,
Stan H.